# Jet transport coefficients by elastic and radiative scatterings in the strongly interacting quark-gluon plasma

### v1

@@ VII Conclusions
We have studied the jet transport coefficients such as $\hat{q}$ – the squared average transverse momentum exchange between the medium and the fast parton per unit length – as well as $\Delta E=dE/dx$ – the energy loss per unit length, including the elastic and inelastic reactions of jet partons with the thermal sQGP described by DQPM in terms of strongly interacting quasiparticles which properties are extracted from the comparison to the lQCD equation-of-state. The transport coefficients have been calculated within kinetic transport theory by accounting for the independent elastic or inelastic collisions of a jet parton with a thermal parton. The contributions of elastic ($q+q\to q+q$ and $q+g\to q+g$) and inelastic ($q+q\to q+q+g$ and $q+g\to q+g+g$) reactions have been calculated by evaluating the leading order Feynman diagrams with effective propagators and vertices from the DQPM by accounting for all channels and their interferences without approximations Grishmanovskii et al. (2023).
We have studied different scenarios for elastic and inelastic scattering of jet parton with a parton from the sQGP medium: we start with the ”default” DQPM calculations where one assumes that all scattered partons, as well as the emitted gluon for inelastic reactions, are in thermal equilibrium. However, since the scattered jet parton and the emitted gluon might be out of equilibrium, we additionally have investigated the influence of different choices for the strong coupling constant $\alpha_{s}$ in thermal, jet parton, and radiative vertices as considered in the literature.
@@
$\bullet$ We have found a strong dependence of transport coefficients on the choice of the strong coupling $g$ used in thermal, jet parton and emitted gluon vertices (cf. Table 1). This dependence is stronger for the inelastic reactions compared to elastic ones due to the extra $g^{2}$ in the gluon emission vertex.
$\bullet$ A strong rise of $\hat{q}$ for elastic and inelastic collisions and an increase of $dE/dx$ for inelastic reactions in the vicinity of $T\to T_{c}$ for the default DQPM scenario is related to the rapid rise of the DQPM coupling $g(T)$ at low $T$ – cf. Fig. 2. The scenario with constant $\alpha_{s}=0.3$ shows a decrease of transport coefficients at low $T$ which is related to the structure of the matrix element squared of the DQPM. Other scenarios with different momentum-dependent jet parton and gluon emitted vertices show a flatting of $\hat{q}$ and a less steep decrease of $dE/dx$ at low $T$.
@@
$\bullet$ We have observed a strong dependence of inelastic $\hat{q}$ on the mass of the emitted gluon – $\hat{q}$ is larger for $M_{g}=0$ compared to the thermal mass of the emitted gluon $M_{g}=M_{g}(T)$ for all jet momenta $p$; $\hat{q}$ decreases with increasing mass of the emitted gluon; similar observations hold for $dE/dx$.
$\bullet$ We have checked the validity of the scaling $\eta/s\approx 1.25\,\frac{T^{3}}{\hat{q}}$ proposed in Ref. Müller (2021) for the DQPM model. We have demonstrated that the ratio $T^{3}/\hat{q}$ to $\eta/s$ strongly depends on the choice of $\alpha_{s}$ in the scattering vertices. It rises steeply with a lowering of $T$ for the case of a temperature-dependent strong coupling $g(T)$, only at large $T$ the ratio falls to the predicted scaling value of 1.25 Müller (2021), while for constant $\alpha_{s}$ and transverse momentum dependent $g$ the ratio approaches a constant already at $T~{}1.5T_{c}$. Thus, in line with the findings in Ref. Karmakar et al. (2023) for the HTL-based model, this scaling is valid only in the weak coupling regime of large $T$ and violated for the strong coupling regime of small $T$.

### v2

@@ VII Conclusions
We have studied the jet transport coefficients, such as $\hat{q}$ – the squared average transverse momentum exchange between the medium and the fast parton per unit length – as well as $dE/dx$ – the energy loss per unit length, including the elastic and inelastic reactions of jet partons with the thermal sQGP described by DQPM in terms of strongly interacting quasiparticles whose properties are extracted from the comparison to the lQCD equation of state. The transport coefficients have been calculated within kinetic transport theory by accounting for the independent elastic and inelastic collisions of a jet parton with a thermal parton. The contributions of elastic ($q+q\to q+q$ and $q+g\to q+g$) and inelastic ($q+q\to q+q+g$ and $q+g\to q+g+g$) reactions have been calculated by evaluating the leading-order Feynman diagrams with effective propagators and vertices from the DQPM, accounting for all channels and their interferences without approximations Grishmanovskii et al. (2024).
We have studied different scenarios for elastic and inelastic scattering of jet parton with a parton from the sQGP medium, starting with the ”default” DQPM calculations, where one assumes that all scattered partons, as well as the emitted gluon for inelastic reactions, are in thermal equilibrium. However, since the scattered jet parton and the emitted gluon might be out of equilibrium, we have additionally investigated the influence of different choices for the strong coupling constant $\alpha_{s}$ in thermal, jet parton, and radiative vertices, as considered in the literature.
@@
$\bullet$ We have found a strong dependence of transport coefficients on the choice of the strong coupling $g$ used in thermal, jet parton, and emitted gluon vertices (cf. Table 1). This dependence is stronger for the inelastic reactions compared to elastic ones due to the extra $g^{2}$ in the gluon emission vertex.
$\bullet$ A strong rise of $\hat{q}$ for elastic and inelastic collisions and an increase of $dE/dx$ for inelastic reactions in the vicinity of $T\to T_{c}$ for the default DQPM scenario is related to the rapid rise of the DQPM coupling $g(T)$ at low $T$ – cf. Fig. 2. The scenario with constant $\alpha_{s}=0.3$ shows a decrease of transport coefficients at low $T$, which is related to the structure of the squared matrix element of the DQPM. Other scenarios with different momentum-dependent jet parton and gluon emitted vertices show a flatting of $\hat{q}$ and a less steep decrease of $dE/dx$ at low $T$.
@@
$\bullet$ We have observed a strong dependence of inelastic $\hat{q}$ on the mass of the emitted gluon – $\hat{q}$ is larger for the case with $M_{g}=0$ compared to the case with thermal mass of the emitted gluon $M_{g}=M_{g}(T)$ for all jet momenta $p$; $\hat{q}$ decreases with increasing mass of the emitted gluon; similar observations also hold for $dE/dx$.
$\bullet$ We have checked the validity of the scaling $\eta/s\approx 1.25\frac{T^{3}}{\hat{q}}$, proposed in Ref. Müller (2021), for the DQPM model. We have demonstrated that the ratio $\eta/s$ to $T^{3}/\hat{q}$ strongly depends on the choice of $\alpha_{s}$ in the scattering vertices. It rises steeply with a lowering of $T$ for the case of a temperature-dependent strong coupling $g(T)$, only at large $T$ the ratio falls to the predicted scaling value of 1.25 Müller (2021), while for constant $\alpha_{s}$ and transverse momentum dependent $g$ the ratio approaches a constant already at $T\approx 1.5T_{c}$. Thus, in line with the findings in Ref. Karmakar et al. (2023) for the HTL-based model, this scaling is valid only in the weak coupling regime of large $T$ and violated for the strong coupling regime of small $T$.